OpenStudy (anonymous):

Need help in financial applications, please. Joanne and Bill are financing $305,500 to purchase a house. They obtained a 20/6 balloon mortgage at 5.75%. What will their balloon payment be? $206,311.68 $401,524.25 $152,285.77 $249,295.69

OpenStudy (amistre64):

first calculate the monthly payment of a 20 fixed loan at the given rate

OpenStudy (amistre64):

use that to calculate the remaining balance due after 6 years of payments

OpenStudy (amistre64):

I tend to develop my own formula for this:\[A = Bk^n+P\frac{1-k^n}{1-k}\] 20 years of 12 payments a year = n B is the starting balance k is the compounding rate (1+r/12) when A=0 the loan is paid off so we can solve for P, the payments \[0 = Bk^n+P\frac{1-k^n}{1-k}\] \[-Bk^n=P\frac{1-k^n}{1-k}\] \[-Bk^n\frac{1-k}{1-k^n}=P\]

OpenStudy (amistre64):

we can use the same formula after knowing P and let n = 6 years of 12 payments .... to find the remaining balance to pay off

OpenStudy (amistre64):

if we want to do some algebra to it ... \[A = Bk^n+P\frac{1-k^n}{1-k}\] \[A = Bk^n+(-Bk^n\frac{1-k}{1-k^n})\frac{1-k^m}{1-k}\] \[A = Bk^n-Bk^n\frac{1-k^m}{1-k^n}\] \[A = Bk^n(1-\frac{1-k^m}{1-k^n})\] \[A = Bk^n(\frac{1-k^n-1+k^m}{1-k^n})\] \[A = Bk^n(\frac{k^m-k^n}{1-k^n})\] but thats just moving things about

OpenStudy (amistre64):

lol, thats why i just use the first stuff

OpenStudy (amistre64):

what do you get for your monthly payment?

OpenStudy (amistre64):

im getting about 2144.87 for payments and balancing out at about 247k

OpenStudy (amistre64):

if they are using tables and not being as exact ... that 249 seems to be the closest due to rounding errors and such. but thats just a hunch

OpenStudy (anonymous):

you are too fast. I got 2144.51 for monthly

OpenStudy (anonymous):

I got 151894 for the balloon payment. not one of choices

OpenStudy (anonymous):

I also have another formula and got 247096 for the balloon payment. I'm confused.